OpenStudy (zab505):

A number a is a root of P(x) if and only if the remainder, when dividing the polynomial by x - a, equals zero. A. True B. False

OpenStudy (anonymous):

True! You can tell because (x - a) is what we can assume to be a factor...if you solve for a: x - a = 0 + a on both sides x = a THAT is your root. So it will always be the opposite of the number in the parentheses of the factor. You can't always assume something is a factor, but here you can b/c you're looking for a root. Best of luck!

OpenStudy (zab505):

Thank you!